### v4

# AI-based Traffic Modeling for Network Security and Privacy: Challenges Ahead

Abstract
Network traffic analysis using AI (machine learning and deep learning) models made significant progress over the past decades. Traffic analysis addresses various challenging problems in network security, ranging from detection of anomalies and attacks to countering of Internet censorship. AI models are also developed to expose user privacy risks as demonstrated by the research works on fingerprinting of user-visiting websites, IoT devices, and different applications, even when payloads are encrypted.
Despite these advancements, significant challenges remain in the domain of network traffic analysis to effectively secure our networks from evolving threats and attacks. After briefly reviewing the relevant tasks and recent AI models for traffic analysis, we discuss the challenges that lie ahead.
1 Introduction
The sophistication of networks attacks has increased over time due to the rapid proliferation of new technologies, applications, network protocols and devices. Meanwhile, network traffic keeps increasing, both in volume and speed, further complicating security challenges. To effectively secure the networks, traffic analysis has long been recognized as a critical component. As payloads in network packets are hardly available for signature-based detection with the high adoption of TLS by different applications (Google-transparency-mail; Google-transparency-HTTP), statistical models are becoming inevitable to analyze network traffic. Decades of research efforts have led to the development of various statistical, machine learning (ML) and deep learning (DL) models of different capabilities to tackle multiple network security and privacy (NetS&P) tasks.
DL models, in particular, are valuable for network traffic analysis in NetS&P tasks for two key reasons, or rather, to address two significant challenges. First, network data is enormous; for example, a 1 Gbps link generates hundreds of gigabytes of data per day under high utilization, even when considering only packet headers (excluding the payload). And network bandwidth has increased steadily, with even consumer (broadband) bandwidth now reaching 1 Gbps. Meanwhile, enterprises and telcos/ISPs utilize bandwidth capacities that are orders of magnitude higher. Second, the number of features that can be extracted from network traces can potentially reach several hundreds or even thousands in count. For instance, when modeling a user browsing session as a data point for training/inference, a small session can easily consist of 200 packets, with each packet containing 10 attributes (e.g., packet size, inter-arrival time, and so on). Consequently, representing this relatively short session requires 2,000 features. As we know, DL models are able to process and learn useful patterns from huge datasets with large feature space.
In light of the current state-of-the-art of AI in network traffic modeling, this work discusses the key challenges that lie ahead in securing networks and protecting user privacy.
2 NetS&P: Tasks and Models
We discuss important NetS&P tasks and AI-based solutions in the literature, without aiming for exhaustiveness.
2.1 Anomaly detection
Anomalous behaviors occurring on endpoints that manifest in network traffic are the anomalies we aim to detect. In an enterprise network, such anomalies may arise from the adoption of new applications (e.g., ChatGPT), sudden use of new protocols—say, encrypted DNS protocols such as DNS-over-HTTPS (DoH), or DNS-over-TLS (DoT), introduction of new devices (e.g., smartwatches), and so on. At homes or hotel rooms, the presence of hidden IoT devices is a security anomaly (Lumos-hidden-IoT-USENIX-SEC-2022). Malicious activities such as bot communications, malware traffic and DDoS attacks are also anomalies. For network administrators, these anomalies pose potential threats that must be detected and analyzed to determine appropriate mitigation actions, such as implementing new policies for newly introduced devices or blocking IP addresses associated with identified bots.
AI models: The primary challenge for this task stems from the diverse types of network anomalies, including those that are previously unknown. For instance, while the recent encrypted DNS protocols like DoH provide better privacy for users, they also offer malware developers new ways to hide their communications (survey-encrypted-DNS-mw-2022), say, for exfiltrating sensitive information (real-time-DNS-exfiltration-NDSS-2024). These are relatively new anomalies unknown before DoH was deployed. Existing research in literature tackles this challenge of anomaly detection by developing unsupervised and semi-supervised models. Early research works explored PCA (Principal Component Analysis), statistical hypothesis testing and regression models to detect anomalous patterns in network traffic (PCA-Diot-2004; brauckhoff2009applying; NADA-2018; evidence-gathering-2017). With the emergence of generative deep learning models for unsupervised problem settings (AD-robust-AE-2017; adversarially-learned-AD-ICDM-2018), we witnessed the development of reconstruction-based unsupervised network anomaly detection solutions in the last decade (KitSUne-2018-NDSS; GEE-2019). For example, GEE (GEE-2019) trains a variational autoencoder (VAE) on noisy benign network traffic sessions to learn its corresponding representation in the latent space, which it subsequently uses to detect deviating behaviors during the inference stage.
2.2 Attack classification
While anomaly detection solutions are useful for detecting broad and unknown types of threats and attacks, they do not identify the specific attack type, which is important to decide on the appropriate mitigation strategy. For example, DDoS traffic must be immediately blocked as close to the sources as possible (a challenge in itself); whereas if a malware is detected, the infected machine has to be contained, incident response initiated, and forensic analysis triggered. Identifying the type of attack is achieved by classifying the network data into one or more known attack classes of interest, such as botnet activity, DDoS attacks, C&C communications, and password spraying, among others. Anomaly detection and attack classification typically complement each other, e.g., as demonstrated in (ADEPT-2021), enabling i) the detection of both known and unknown attacks, and ii) the identification of attack types.
AI models: While the problem might appear to be a multi-class classification problem, the data available for different attacks vary. DDoS requires simple features that capture the traffic rate, whereas detecting the presence of malware that uses DGA (domain generation algorithm) requires analysis of (plain-text) DNS payload. Therefore, solutions in this space typically train binary classifiers to detect specific attacks. For example, several ML algorithms for DDoS detection are evaluated considering both accuracy and inference latency in (DDoS-eval-latency-ICNP-2024). Similarly, prior works have proposed methods for classifying DGA domains (DGA-Rossow-2024), identifying bot traffic to servers using GAN-based data augmentation (botnet-S&P-2020), and detecting encrypted malware communications over TLS (anderson2017machine) as well as Tor (Tor-malware-CCS-2022).
2.3 IoT device identification
Identifying IoT devices connected to a network is necessary for security auditing and for configuring appropriate policies (e.g., firewall rules) for effective management of the network. The challenge lies in determining the type of devices connected to the network through traffic analysis. In an enterprise network, a device identification solution would analyze north-south traffic as well as east-west traffic to identify devices that communicate both internally and externally. When labels for IoT devices and their corresponding network data are available, the task transforms into a multi-class classification problem.
AI models: IoT device identification has received considerable attention over the past decade, resulting in the proposal of various supervised and semi-supervised approaches for classifying known and unknown devices, respectively. They include several works that are based on conventional ML as well as DL models (2017-profileiot; 2019-DIoT; DEFT-2019; IoT-biLSTM-2020; WU-ZEST). A recent work (usenix-2020-youarewhatyou) demonstrates that devices connected to a WiFi network can be identified by training a DL model on the broadcast/multicast traffic observed.
Device identification is considered a defensive security measure, as organizations (WU-ZEST) and individuals (Lumos-hidden-IoT-USENIX-SEC-2022) need to track the devices connected to their networks. However, when viewed from the perspective of a home or public user, adversarial identification of consumer devices—whether using a compromised router or through a public WiFi hotspot—represents a violation of user privacy. This has led to the development of counter-fingerprinting techniques, e.g., using generative adversarial perturbations (shenoi2023ipet).
2.4 Website fingerprinting (WFP) attack
WFP attack targets user privacy by aiming to identify the websites a user visits. The research efforts in this space shed light on the vulnerability of existing and new network protocols in revealing sensitive user information. Since the focus is on web traffic, the relevant network protocols for traffic modeling are primarily limited to DNS (and its encrypted variants) and HTTPS. However, as users may employ VPNs or Tor (dingledine2004tor) to mitigate such privacy attacks, research proposals typically assume that the traffic is tunneled through these services. The targets of WFP attacks can include anyone, such as citizens of interest, journalists, activists, or politicians, who may be monitored (or surveilled) by those in power. Consequently, the threat model assumes an adversary that has access to a network node, such as a router (either through direct control or compromise). Thus the adversary can passively monitor and analyze network traffic to identify the websites visited by the target users.
AI models: Given that a browsing session consists of hundreds or even thousands of packets, this topic has witnessed significant development of deep learning-based attack strategies. In other words, the ability of deep learning models to learn from a large feature space represents a substantial advantage for this task. The literature has interesting WFP attack proposals that use various DL architectures ranging from CNN to AE to sequence models such as LSTM and transformer (NDSS-2018-wfp-lstm; pfp-PETS-2019; varCNN-WFP-2019; Usenix-2022-onlinewebsitefingerprinting; webpage_quic; multi-tab-FP-transformer-SP2023; transformer-WFP-CCS-2023; MIL-WFP-USENIX-2024; Levi-2024). There are also strategies proposed for countering them (gong2020FRONTGLUE; smith2022QCSD; GAN-WFP-defense-SP-2022; siby2023PET; Usenix-2023-website-fingerprinting). These models are evaluated under two scenarios: i) a closed-world (lab) setting, where a classification model is trained to classify $n$ specific monitored websites (each with samples collected over time) and is tested for accurate classification of the same $n$ websites, where $n$ is typically in a few hundreds; ii) a more realistic open-world setting, where the challenge is to identify $n$ monitored websites under the assumption that the user visits $m~(m\gg n)$ unmonitored websites, where $m$ is in tens of thousands. Since tunneling of traffic hides packet headers, (different from the previous tasks) the features used here are the basic (or raw) ones such as packet-size, inter-arrival time (IAT) (or time lag) of packet and direction of packet (as browsing generates bidirectional traffic).
2.5 Other tasks and models
There are a numerous other NetS&P tasks; due to space constraints, we provide a brief outline here.
  • Censorship and Anonymity. Nation-state adversaries carry out traffic analysis for surveillance and censorship. The Tor anonymity network (dingledine2004tor) is developed to offer anonymized access to Internet users. However, Tor is prone to de-anonymization attack (nasr2018deepcorr; oh2022deepcoffea). A most recent proposal (RECTor-2025) slices flows into windows and employs attention-based MIL (multi-instance learning) to learn the most relevant segments of potentiality noise traffic flows, so as to correlated flows at the entry and exit nodes of a Tor network.
  • Token inference attack. In this attack, the encrypted traffic is passively monitored to estimate the token sizes from response packets, to subsequently use to to infer the response from an AI assistant (prompt-infer-USENIX-SEC-2024), using a fine-tuned T5 transformer model (T5) is fine-tuned.
We also highlight the application of adversarial ML, which aims to generate adversarial samples for evading systems utilizing AI models. In the context of NetS&P, an adversarial ML approach can be exploited not only to bypass AI-based security systems but also to counter threats related to privacy and censorship. For instance, authors in (shenoi2023ipet) adapt generative adversarial perturbations (poursaeed2018generative) for evading IoT fingerprinting solutions that employ AI models. And in the space of website fingerprinting, a defense based on GAN, specifically, WGAN-div (wu2018wasserstein), to evade WFP attacker (discriminator) is studied in (GAN-WFP-defense-SP-2022).
3 Challenges (and Opportunities) Ahead
3.1 Data challenges
One of the most important questions is, do the current AI-based solutions generalize beyond the experiments conducted in controlled lab environments? Addressing this question is challenging due to the lack of high-quality labeled data. Most research on network anomaly detection, attack classification, and device identification relies on openly available labeled datasets. However, since these datasets are generated in lab environments, they often contain artifacts or ‘bad design smells’ (bad-smells-EuroSP-2024) that raise concerns on model over-fitting and biases. In  (bad-smells-EuroSP-2024), the authors analyze seven highly-cited datasets and provide evidence of such artifacts in network intrusion detection research. For instance, in one dataset with eight attack classes, two basic features exhibit minimal overlap between benign and attack traffic. As a result, a simple perturbation of these two features—unrelated to the inherent characteristics of attacks—can evade multiple ML and DL models. This simple approach outperforms a state-of-the-art adversarial attack (adversarial-NIDS-2022) employing evolutionary computation and generative adversarial networks. This finding suggests that evaluations in (adversarial-NIDS-2022) may not provide a “meaningful measure of the attack’s effectiveness” (bad-smells-EuroSP-2024) due to dataset flaws. Similar biases in network datasets have also been highlighted for other NetS&P tasks, e.g., website fingerprinting (data-explainable-WFP-PETS-2023; retrace-Tor-2024; genuineTorTraces-2024).
The issue at hand is not about the existing proposals per se, but rather the quality of the datasets used for evaluating network security tasks. Going one step further, the fundamental challenge is the difficulty of validating the quality network traffic. The traffic generated by a single day’s communications in an enterprise network can reach hundreds of gigabytes, and when considering this scale over multiple days, across various enterprises and different verticals, the volume becomes staggering. Therefore, collecting and labeling real-life network traffic flows is an obstacle. Furthermore, real-world network traffic contains sensitive user information, and sharing this data poses significant privacy risks and ethical concerns. Model-based attacks can extract hidden patterns as fingerprinting attacks in the previous section reveal; meta-data from packet headers (excluding payloads) is sufficient to infer sensitive information. Consequently, it is understandable that enterprises and telco operators refrain from publicly sharing traffic collected from their networks for research and development purposes.
There is currently a lack of real-world quality datasets with labels for NetS&P tasks such as anomaly detection, attack detection, website fingerprinting, etc.
This leaves the research community with the second-best option—generating traffic datasets in controlled lab environments. There are promising directions here. The aforementioned in-depth analysis of open datasets represents an important step forward; it also sheds light on how to evaluate the quality and fidelity of lab-generated datasets. Learning of broad principles will guide in framing the right questions for specific NetS&P tasks. For example, if SSH brute-force attempts are present in the attack category, the benign category should also have normal SSH flows; otherwise encoding the port (22) would classify all SSH flows as malicious and (falsely) validate the model!
We need to establish design principles for generating real traffic data in controlled environments, and the data generated should be evaluated for fidelity.
As second direction, we note several recent research proposals put forth to synthesize traffic data with high fidelity (botnet-S&P-2020; GAN-pkt-synthesis-SIGCOMM-2022; data-synthesis-DCN-HotNets-2023; Zoom2Net-SIGCOMM-2024; jiang2024netdiffusion; flowChronicle-2024). In (data-synthesis-DCN-HotNets-2023), the authors propose using a Variational Autoencoder (VAE) to learn the flow-size distribution from datacenter network traces, subsequently generating sequences of flow sizes with a Recurrent Neural Network (RNN) using Gated Recurrent Units (GRU). In (Zoom2Net-SIGCOMM-2024), a transformer model is used to estimate missing data in a network telemetry time-series data by correlating information from multiple sources. Although they do not directly address security problems, we can take a leaf out of them to understand the key requirements that have to be met to synthesize data for NetS&P tasks:
Synthetic Data Generation: How critical is the leakage of distribution parameters in the context of enterprise network data? How can models learn with minimal linkage of distribution parameters to the source network? What impact does constrained data synthesis have on performance of NetS&P solutions?
Another promising direction is emulating applications that generate network traffic in a controlled environment. This is a common practice in website fingerprinting domain, where thousands of websites are visited (over Tor and VPN) using browsers, to generate the corresponding browsing traffic (e.g., HTTPS packets). However, for training models for securing a consumer or enterprise network, broader set of applications have to be emulated. A recent proposal explores the possibility of generating network traffic by orchestrating public GitHub repositories (gen-nw-traffic-2022). Building on this concept, the authors of  (khan2024harnessing) propose a data-generation platform that can be managed using SDN. We are still in early stages, and further research is needed to validate the traffic thus generated, and subsequently also to explore generating malware and anomalous traffic with high fidelity and diversity.
Emulation of benign and malicious traffic: How do we emulate applications corresponding to different enterprises (e.g., finance vs. academia) with varying user characteristics, for generating benign application data? How do we carefully and ethically emulate malware behavior, including infection, lateral movement, C&C communications, data exfiltration, and other activities, for the purpose of generating malicious data? How can SDN help to dynamically configure and manage networks for traffic generation? (netreplica-2025)
3.2 Practical deployment
Real-time per-packet inference in network traffic analysis faces several practical limitations. Enterprise networks operate at tens of Gbps whereas telco networks operate at 100s of Gpbs. Even at 10 Gbps, the time available for per-packet decision-making is constrained to less than 100 nanoseconds per packet, which poses challenges for complex computations required for DL-based solutions. In fact, an expensive step in building an ML/DL pipeline is parsing packets for high-speed feature extraction (Unsenix2023-IDP-programmable-switch-flow).
A single packet typically does not provide sufficient information to make informed decisions about its legitimacy. For instance, a TCP SYN packet can be part of a legitimate connection or a malicious attempt.Only by aggregating packets, e.g., by source/destination IP address/port and segmented by time, can patterns indicative of attacks, such as TCP SYN floods, be identified. However, storing and processing these packet aggregates require substantial resources, a challenge that increases with the network rate.
Packet aggregates serve as meaningful units for modeling traffic. However, extracting information—such as total volume or duration—from packet aggregates (e.g., 5-tuple flows or sessions grouped by src/dst IP) still necessitates processing each individual packet, which becomes increasingly challenging as network rates rise.
In this context, programmable data planes offer promising solutions for in-network computation at rates of terabits per second (Tpbs) (in-network-dumb-switch-hotnets-2017; zhang2020poseidon; Usenix-DDos-programme-switch-2021; p4-survey-2021). These are essentially switches and smartNICs that allow programmability of packet processing pipelines directly on the network devices while constraining to specific packet processing behavior, e.g., in terms of latency and throughput. Indeed, there have been attempts to implement tree-based models in programmable switches as that aligns well with the match-action pipeline of the programmable switch architecture (2019-do-switchdream; Unsenix2023-IDP-programmable-switch-flow; offloadingML-ACM-Surveys-2023; hybrid-inetwork-classification-2024).
However, the cost of processing every single packet, even when decisions are based on packet aggregates, remains a significant challenge (Marina-ML-monitoring-Terabit-2024). Sampling has long been a viable solution for various network tasks, such as heavy-hitter monitoring (sampling-elephants-2003). Nonetheless, applying sampling to model network traffic for security tasks requires further investigation. Developing models that can effectively learn from and infer dynamic features is important. Additionally, we have to design intelligent sampling strategies tailored to security use cases. For instance, protocol handshakes (TCP, TLS, etc.) are carried out in the initial part of a connection, and therefore might reveal important information (C2-TLS-RAID-2024). This raises a research question—would an adapting sampling strategy that samples different parts of a packet sequence at different probabilities be better than static sampling methods (IPFIX)? To answer this, we need to explore the trade-offs between sampling methods and their impact on model performance in detecting security threats. With programmable data planes, we now have the possibility to develop an in-network solution that makes both decisions—sampling and inference—without leaving the data plane.
Programmable data planes hold promise for implementing in-network DL-based solutions. Intelligent sampling, combined with models that can effectively handle missing data points, need to explored for NetS&P tasks.
3.3 Explainability of predictions
Explainability technically allows us to interpret the behavior of models, or in other words, decisions made by models. The capability to explain the model predictions is important in cybersecurity to build trust between model developer, end user and stakeholders (XAI-depoloyment-2020). Explainable models also aid in improving performance by identifying and mitigating potential pitfalls, such as biases, vulnerabilities to out-of-sample data, and overfitting (Trustee-CCS-2022; bad-smells-EuroSP-2024).
Consider different cybersecurity use cases. For instance, in malware analysis, an explainable model highlights specific API calls that perform malicious actions or endpoint registers that were modified, providing actionable insights for mitigation (XAI-malware-2024). In phishing detection, an explanation-based solution can reveal discrepancies between the brand displayed on a webpage and its domain name, helping users understand why a website is flagged as suspicious (lee-LLM-phishing-2024). For provenance graph modeling of endpoint events (prov-ninja-2023), a SOC (Security Operations Center) analyst would find it helpful if a model identifies the sequence of processes and events that led to the alert raised, such as executing a dropped binary that connects to a malicious IP.
The above examples illustrate the utility of local explanations that explain (the features leading to) inference decision for the given input samples. On the other hand, with global explanations, the goal is to decipher the overall understanding of the model; e.g., knowing certain expensive features do not increase the detection capability is useful to reduce the cost of deployment (cost-aware-FS-IoT-2021; Marina-ML-monitoring-Terabit-2024). We refer readers to (zhao2024explainability) for a detailed breakdown of these two categories of explanations and understanding of the current state-of-the-art techniques for explainability. Some of these techniques have been recently explored in the network security community, concentrating mostly on anomaly and attack detection (GEE-2019; Trustee-CCS-2022; xNIDS-USENIX-SEC-2023; global-explanations-AD-CCS-2024). The state-of-the-art DL solutions for NetS&P tasks, e.g., (multi-tab-FP-transformer-SP2023; WU-ZEST; prompt-infer-USENIX-SEC-2024; RECTor-2025), are primarily based on the transformer model architecture (vaswani2017attention); besides, transformer-based foundation models are also emerging for NetS&P tasks (see below). This indicates that the existing techniques and challenges related to explainability in transformer models are pertinent to NetS&P.
We also have to take consider the perspective of the end-user to understand what needs to be ‘explained’ along with the inference results. And this is task dependent. In anomaly or attack detection, a SOC analyst needs to understand why a particular traffic session is deemed malicious. This requires going beyond ranking important features (used for input representation) (GEE-2019; Trustee-CCS-2022). Indeed the prediction should be accompanied with concrete explanations, such as noting that ‘there is a series of connection attempts from this enterprise endpoint that failed,’ indicating a C&C connection attempt from an infected endpoint. A potential approach might be to utilize LLMs to ‘translate’ input feature-level explanations to a format consumable by SOC analysts.
Explainability also provides insights into adversarial ML attack capability (and thereby defense); e.g., DGA classifiers are known to be prone to adversarial attacks (DGA-Rossow-2024). In counter-censorship models, explaining the results would aid in understanding the censorship strategy used (detecting-proxies-NDSS-2020; shadowsocks-China-IMC-2020; VPN-fingeprinting-2025)—such as whether the censorship technique in use depends highly on the features of the first few packets (shadowsocks-China-IMC-2020)—so as to build effective models for evading censorship. Knowing which features contribute to the accuracy of a model (iglesias2015analysis; cost-aware-FS-IoT-2021) is useful in estimating the cost of deploying a solution.
Understanding task-specific user requirement is important to determine the explainability capability we expect from a model. User studies (SoK-NIDS-assessment-2023) help to gain a better understanding of requirements from stakeholders on explainable models for specific use cases.
3.4 Many models; convergence on the horizon
For more than two decades, network traffic modeling progressed alongside the advancements in AI, resulting in the development of hundreds of models for NetS&P tasks. We moved from simple statistical methods to ML classifiers such as Decision Trees to more complex and powerful DL models such as transformers; this evolution has been instrumental in training models with bigger datasets and larger feature space, allowing researchers to also define numerous features for various tasks.
Given that network protocols are clearly defined, and substantial progress has been made over the past decades in developing AI-based solutions for NetS&P tasks, it can be argued that the fundamental features useful for modeling are well understood. For illustration, consider a set of research works across the aforementioned NetS&P tasks—anomaly detection (KitSUne-2018-NDSS; GEE-2019), flow classification for botnet detection (FlowLens-NDSS-2021), muti-class attack classification on IoT devices (ADEPT-2021), multi-tab website fingerprinting (multi-tab-FP-transformer-SP2023), flow-correlation attack on Tor networks (oh2022deepcoffea), video fingerprinting (schuster2017beauty; sabzi2024netshaper), dApp fingerprinting (dapp-GNN-FP-2021), prompt inference (prompt-infer-USENIX-SEC-2024) and Tor de-anonymization attack (RECTor-2025)—the raw or fundamental packet-level features at the finest granularity used are packet-size, inter-arrival time (IAT) between packets, direction, protocol, and representations for src/dst IP addresses and port numbers. Features at a coarser granularity, such as flow-level features, are derived from these raw features; for example, flow volume (duration) is the sum of sizes (IAT) of packets in that flow [^1] To be sure, there are a few exceptions. For example, DGA detection is based on DNS payloads, which however would not be available when encrypted DNS protocols are employed. Another exception is TLS feature extraction for representing session-level information. However, TLS is ubiquitous, and the transition from TLS 1.2 to TLS 1.3 has made most features unavailable, thanks to the encryption of handshakes and introduction of ECH. The estimation of certificate size that might be useful for C&C detection (C2-TLS-RAID-2024), could be learned by training on packet-sequences with size being one of the features..
State-of-the-art: Fundamental features from raw encrypted packets remain largely consistent across various NetS&P tasks. DL models today are capable of handling huge datasets and large feature space. A natural next challenge is to go beyond having multiple models for the different tasks and build a foundation model for network traffic analysis.
The concept of foundation models recently gained attention in the network community (foundation-model-thinking-HotNets-2022; ET-BERT-2022; Ptu-pre-trained-nw-traffic-ICNP-2024; llm-cybersec-2024; netfound-2025). Once a foundation model has been built, it can be applied to different NetS&P tasks, say, by fine-tuning with task-specific data. We argue that, for a foundation network model to be truly effective, it must possess several key properties that enhance its practical utility:
  • Different network environments—consumer, enterprise, datacenter, and telco (5G network)—have different capacities and characteristics, leading to distinct methods of traffic capture. Cost of extracting and storing traffic features is a primary constraint, influenced by the speed and scale of the network—1 Gbps consumer bandwidth with a few nodes compared to a telco network operating at hundreds of Gbps with millions of subscribers. Therefore, a flexible foundation model is necessary to make inferences on traffic captured at different scales represented at multiple levels: i) raw packet features, ii) flows or 5-tuple packet aggregates, and iii) sessions or flow aggregates (UniNet-2025). This flexibility enables the model to be deployed across networks with diverse data rates and capabilities. For instance, a home consumer network might have the capability to extract raw features from all packets, thereby deriving flow and session features, and thus making all views available. In contrast, a telco network, handling much higher volumes of traffic, might only collect aggregate features, such as IP-based session-level data.
  • Following the unsupervised pretraining process in NLP (BERT-2018) (e.g., using masked language modeling), foundation network model could be trained in a self-supervised way. However, fine-tuning to specific use cases requires labeled data. Labeling flows and packets is not only labor-intensive but also prone to errors. Therefore, a foundational model should be designed to handle missing labels during fine-tuning for specific tasks, enabling it to adapt effectively even when complete annotations are not available.
  • As multiple foundation models are likely to emerge out of research, each potentially trained on different datasets—some openly available (e.g., MAWI (MAWI-dataset) and CAIDA (caida-2018-2019) and others private (ET-BERT-2022; Ptu-pre-trained-nw-traffic-ICNP-2024)–it becomes important to explore methods for combining these models effectively for better cost-effective utilization. Simultaneously, we must address the challenge of minimizing the risk of sensitive information leakage when integrating these models, ensuring that the benefits of collaboration are realized while maintaining data privacy and security.
  • Explainability is an active and challenging issue in foundation models today (zhao2024explainability). When designing a foundation model for NetS&P tasks, it is crucial to prioritize explainability. The challenge lies in utilizing the foundation model for various downstream tasks while also providing clear interpretations of its decisions.
Key properties of a foundation model for network traffic: Representations capturing different traffic characteristics at different scales and views. Capability to handle missing and noisy labels. Collaboration and openness, while ensuring user privacy is protected, for the effective utilization of multiple foundation models. Explainability for the downstream tasks.
4 Conclusion
In this work, we highlighted the advancements in traffic modeling for NetS&P tasks, with the aim of understanding and identifying the challenges that lie ahead. We hope this discussion serves as a catalyst for defining relevant research problems and generating new ideas, ultimately enhancing the likelihood of developing deployable and effective models for securing networks and users.
5 Acknowledgment
This research/project is supported by the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under the National Cybersecurity R&D Programme and the CyberSG R&D Programme Office (Award CRPO-GC2-ASTAR-001). Any opinions, findings, conclusions, or recommendations expressed in these materials are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, the Cyber Security Agency of Singapore, or the CyberSG R&D Programme Office.